OpenStudy (anonymous):

Let P(x) + ax^7 + bx^3 +cx -5 where a,b,c are constants. Given that P(-7) = 7, find the value of P(7).

OpenStudy (radar):

What course are you taking, this is very perplexing to me at least. Hopefully robtobey can solve, it is above my paygrade.

OpenStudy (anonymous):

\[ P(x) =ax^7 + bx^3 +cx -5\]?

OpenStudy (anonymous):

note that all the exponents are odd numbers

OpenStudy (anonymous):

\[-P(7)=-a7^7-b7^3-7c+5=a(-7)^7+b(-7)^3+c(-7)-5+10=P(-7)+10\]does that help?

OpenStudy (anonymous):

this is Algebra Grade 9

OpenStudy (anonymous):

This helps mulishly

OpenStudy (anonymous):

very well :)

OpenStudy (anonymous):

what that word means? mulishly?

OpenStudy (anonymous):

sorry..I typed your name and the computer recognized it differently . when I clicked "return" it sent that name...mukushla

OpenStudy (anonymous):

oh...np..:)